OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. M(0, 6), N(6, 0)

OpenStudy (anonymous):

y=-x+6

OpenStudy (anonymous):

Thank you helper michael

OpenStudy (anonymous):

slope is -1 for equation use y-6=-1*(x-0_ y=-x+6

OpenStudy (anonymous):

My problem starts with x